lens maker formula

Grade:12

2 Answers

Chetan Mandayam Nayakar
11 years ago

1/f =(n-1)((1/r1)-(1/r2))

ABHINAV MISHRA
11 years ago

1/f= (1/v-1/u)(n2/n1-1)

where f is focal length 

v and u are image and object distance

and n is refractive index